# Imprecise Markov Semigroups and their Ergodicity

### v4

@@ 4. Ergodicity in the General Case
  • $\overline{\mu}$ and $\underline{\mu}$ are measures on $\mathcal{F}$.
@@
  • The Markov semigroups $\overline{\mathbf{P}}=(\overline{P}_{t})_{t\geq 0}$ and $\underline{\mathbf{P}}=(\underline{P}_{t})_{t\geq 0}$ associated with $\overline{\Gamma}$ and $\underline{\Gamma}$ are the greatest and least elements of an imprecise Markov semigroup $\mathbf{Q}$, [^16] With respect to $\preceq^{\text{part}}_{\tilde{f}}$ or $\preceq^{\text{tot}}_{\tilde{f}}$. respectively.
We are now ready for the main result of this section. It hinges on the concepts of adjoint operator, self-adjointness, essential self-adjointness (ESA), extended algebra $\mathcal{A}$ of the algebra $\mathcal{A}_{0}$ associated with a diffusion Markov triple, connexity, and weak hypo-ellipticity. They are brushed off in B.2. As its proof technique is similar to that of Theorem 8 (despite exploiting different, more general results), we defer it to A.
@@
The connexity and weak hypo-ellipticity assumptions allow to “import” some of the geometric structure of the Riemannian manifolds considered in Section 3 to the more general diffusion Markov triple setting, and to relax the requirement in Theorem 8 that $\underline{L}$ and $\overline{L}$ are elliptic operators. This is expanded upon in B.2.
Similarly to what we pointed out in Remark 1, if we do not know whether the (reversible invariant) measures $\overline{\mu}$ and $\underline{\mu}$ are finite, then in the statement of Theorem 10 we must require that $\rho_{\overline{L}},\rho_{\underline{L}}>0$ (Bakry et al., 2014, Theorem 3.3.23). [^17] So that (11) holds.
Similarly to Corollary 8.1, the following is an important result that subsumes the ergodic behavior of the elements of $\mathbf{Q}$ when applied to $\tilde{f}\in B(E)$.
Corollary 10.1 (Ergodicity of Imprecise Markov Semigroup $\mathbf{Q}$, General).

### v5

@@ 4 Ergodicity in the General Case – Compact $\mathbf{Q}$
  • $\overline{\mu}$ and $\underline{\mu}$ are ($\sigma$-finite) measures on $\mathcal{F}$.
@@
  • The Markov semigroups $\overline{\mathbf{P}}=(\overline{P}_{t})_{t\geq 0}$ and $\underline{\mathbf{P}}=(\underline{P}_{t})_{t\geq 0}$ associated with $\overline{\Gamma}$ and $\underline{\Gamma}$ are the greatest and least elements of a compact imprecise Markov semigroup $\mathbf{Q}$, [^13] With respect to $\preceq^{\text{part}}_{\tilde{f}}$ or $\preceq^{\text{tot}}_{\tilde{f}}$. respectively.
We are now ready for the main result of this section. It hinges on the concepts of adjoint operator, self-adjointness, essential self-adjointness (ESA), extended algebra $\mathcal{A}$ of the algebra $\mathcal{A}_{0}$ associated with a diffusion Markov triple, connexity, and weak hypo-ellipticity. They are brushed off in Appendix C.2. As its proof technique is similar to that of Theorem 10 (despite exploiting different, more general results), we defer it to Appendix A.
@@
The connexity and weak hypo-ellipticity assumptions allow to “import” some of the geometric structure of the Riemannian manifolds considered in Section 3 to the more general diffusion Markov triple setting, and to relax the requirement in Theorem 10 that $\underline{L}$ and $\overline{L}$ are elliptic operators. This is expanded upon in Appendix C.2.
Similarly to what we pointed out in Remark 4, if we do not know whether the (reversible invariant) measures $\overline{\mu}$ and $\underline{\mu}$ are finite, then in the statement of Theorem 12 we must require that $\rho_{\overline{L}},\rho_{\underline{L}}>0$ (Bakry et al., 2014, Theorem 3.3.23).
Similarly to Corollary 10.1, the following is an important result that subsumes the ergodic behavior of the elements of a compact IMSG $\mathbf{Q}$ when applied to $\tilde{f}\in B(E)$.
Corollary 12.1 (Ergodicity of Compact Imprecise Markov Semigroup $\mathbf{Q}$, General).